OpenStudy (anonymous):

ind the discriminant of the function f(x) = x2 - 4x + 5 and determine how many roots the function has.

OpenStudy (maheshmeghwal9):

u can take discriminant as \[b^2-4ac\]=\[(-4)^2-4*1*5=16-20=-4\]

OpenStudy (campbell_st):

the discriminant is \[b^2 - 4ac\] in your question a = 1, b = -4, c = 5 substitute if discriminant > 0 there are 2 unequal roots... and if the discriminant is a perfect square the roots are rational discriminant = 0 there is 1 repeated root. discriminant < 0 there are complex roots..... sometimes called undefined roots

OpenStudy (anonymous):

A) -36; 1 real root B) -36; 2 imaginary roots C) -4; 1 real root D) -4; 2 imaginary roots not c

OpenStudy (campbell_st):

if the value is -4 there are 2 imaginary

OpenStudy (anonymous):

I just would like to expand on campbell_st's post by explaining why this is so... Recall that the quadratic formula states that: \[(-b \pm \sqrt{b^2 - 4ac})/2a\] The discriminant is the expression inside the root, namely \[b^2 - 4ac\] If discriminant > 0, then we get: \[(-b \pm \sqrt{discriminant} )/2a\] Which means the roots are:\[(-b + \sqrt{discriminant} )/2a\] and \[(-b - \sqrt{discriminant} )/2a\] If discriminant = 0, then we get one root, which is: \[(-b)/2a\] //because the root of 0 is 0. If the discriminant is < 0, then we get complex roots because the root of negative numbers yields complex numbers. \[\sqrt{-1} =i\] ^ i is a complex number